OpenStudy (anonymous):

Will give out medals. John kept track of how many baskets were made in a basketball game. After 4 minutes, 5 baskets were made. How many baskets were made after 13 minutes?

OpenStudy (anonymous):

Is there any answer choices with this question, if there is I would like to look at them

OpenStudy (anonymous):

Well, 4 into 13 minutes: 4, 8, 12 = 5 *3 = 15 4 minutes equals 5 baskets 8 minutes equals 10 baskets 12 minutes equals 15 baskets Alright so now you have to calculate how 5 baskets are made in 4 minutes. I would have to say that, if there is any answer choices don't pick the answers below 15, if it's below 15 then it's wrong. Now with no answer choices provided I would have to guess and say 16 or 17 baskets were made in 13 minutes! :p

OpenStudy (anonymous):

Hello? U still here @Sameerah11 ?

OpenStudy (anonymous):

5 baskets in 4 minutes. To find how many/minute, you dived. 5/4 = 1.25 per minute. Now you simply multiply 1.25 by 13. That is your answer.